Grade 11Discuss with Askiitians Tutors

Find the equation of the straight line upon which length of the perpendicular from origin is 2 and the slope of this perpendicular is 5/12??

Miljot Singh
10 Years agoGrade 11

1 Answer

ADITI SINGH
9 Years ago
Arrange your equation into slope-intercept form: 
4x + 3y = 12 
3y = -4x + 12 
y = -4/3x + 4 

Our new line will take the same form of "y = mx + b". First recognize that a perpendicular line has an inverse slope to its reference. 
m(reference) = -4/3 
m(perpendicular) = 3/4 

To solve for "b" we'll plug in (0, 0) since we know the line passes through the origin. 
y = 3/4 x + b 
0 = 0 + b 
b = 0 

Our equation is thus: 
y = 3/4 x 

*Edit* 
I'm sorry, I didn't actually answer your question. These two lines intersect at: 
3/4x = -4/3x + 4 
9x = -16x + 48 
25x = 48 
x = 1.92 

And: 
y = 3/4 x 
y = 3/4 (1.92) 
y = 1.44 

We're finding the hypotenuse length given the above "y" and "x" values, using the Pythagorean Theorem. 
h² = x² + y² 
h² = 1.92² + 1.44² 
h = 2.4 

The length of the requested line is 2.4 (units).